OpenStudy (anonymous):

Verify the identity. Sin (x+y)-sin (x-y)=2 cos x sin y

OpenStudy (mertsj):

\[\sin (x+y)=\sin x \cos y+\cos x \sin y\]

OpenStudy (mertsj):

\[\sin (x-y)=\sin x \cos y-\cos x \sin y\]

OpenStudy (anonymous):

Wait what did you do?

OpenStudy (mertsj):

I gave you two identities that you can use in your problem

OpenStudy (anonymous):

So like I would equal the problem out to zero and then plug the identity?

OpenStudy (anonymous):

so like no like, you would like, work out one side of the equation, by like, subbing in the identities he gave you, like, until like both sides are equal, like.

OpenStudy (anonymous):

Thanks!

OpenStudy (mertsj):

yw